OpenStudy (anonymous):

three identical positive charges Q are arranged at the vertices of an equilateral triangle..the side of the triangle is a..find the intensity of the field at the vertex of a regular tetrahedran of which the triangle is the base...

OpenStudy (anonymous):

 since the charge distribution is symmetric about the axis the net E field is \[E(total)=3Ecos30\] \[=3(\sqrt{3}/2)*(1/4\Pi \epsilon0)*q/a^2\] \[=[27\sqrt{3}/2]*10^{9}q/a ^{2}\]